### v1

# Enhancing Automatic Chord Recognition via Pseudo-Labeling and Knowledge Distillation

Abstract
Automatic Chord Recognition (ACR) is constrained by the scarcity of aligned chord labels, as well-aligned annotations are costly to acquire. At the same time, open-weight pre-trained models are currently more accessible than their proprietary training data. In this work, we present a two-stage training pipeline that leverages pre-trained models together with unlabeled audio. The proposed method decouples training into two stages. In the first stage, we use a pre-trained BTC model as a teacher to generate pseudo-labels for over 1,000 hours of diverse unlabeled audio and train a student model solely on these pseudo-labels. In the second stage, the student is continually trained on ground-truth labels as they become available, with selective knowledge distillation (KD) from the teacher applied as a regularizer to prevent catastrophic forgetting of the representations learned in the first stage. In our experiments, two models (BTC, 2E1D) were used as students. In stage 1, using only pseudo-labels, the BTC student achieves over 98% of the teacher’s performance, while the 2E1D model achieves about 96% across seven standard mir_eval metrics. After a single training run for both students in stage 2, the resulting BTC student model surpasses the traditional supervised learning baseline by 2.5% and the original pre-trained teacher model by 1.55% on average across all metrics. And the resulting 2E1D student model improves from the traditional supervised learning baseline by 3.79% on average and achieves almost the same performance as the teacher. Both cases show the large gains on rare chord qualities.
1 Introduction
Automatic Chord Recognition (ACR) is a fundamental task in Music Information Retrieval (MIR) that aims to identify the harmonic content of audio recordings by outputting a sequence of chord labels. While large-scale labeled datasets are readily available for many machine learning domains, ACR faces significant data constraints: publicly available labeled chord datasets remain limited in both size and diversity [13]. This scarcity stems from the substantial manual effort required for precise audio-label alignment and consistent harmonic interpretation, as chord boundaries are inherently ambiguous and context-dependent [9, 24]. Furthermore, chord vocabularies are large and highly imbalanced; models excel on frequent major/minor chords but underperform on rare seventh and extended chords [13, 24, 1]. By training models only on scarce labeled data, we limit ourselves from making use of the unlabeled audio available in much larger quantities.
Hence, we use unlabeled audio to improve the training of chord recognition and show that the accuracy gains are disproportionately concentrated on rare chord qualities. For this purpose, we propose a two-stage training pipeline (see Figure 3). In the first stage, a pre-trained teacher model generates pseudo-labels for over 1,000 hours of diverse unlabeled audio, and a student model is trained solely on these pseudo-labels until convergence - without requiring any ground-truth annotations. In the second stage, when labeled data becomes available, the pseudo-label-trained student is continually trained on ground-truth labels. To prevent catastrophic forgetting [7] of the representations acquired in the first stage, we apply selective knowledge distillation (KD) from the teacher as a regularizer throughout the second stage.
Prior semi-supervised ACR methods [1, 16] require ground-truth labels from the outset and fuse pseudo-labeled and labeled data together in a single training. Our two-stage training approach decouples labeled and unlabeled data, making the training accessible even when labeled data is initially unavailable. Notably, open-weight pre-trained models are often more readily available than their training data, offering a practical way to leverage teacher knowledge without access to proprietary datasets.
We also introduce a compact, purely transformer-based dual-encoder architecture (2E1D) that is based on [32] and is more lightweight than the teacher model. We demonstrate that our method can generalize across architectures. Our experiments show that the best resulting student surpasses both the traditional supervised learning training method and the pre-trained teacher model across all seven standard mir_eval metrics [26], with particularly large gains on rare chord qualities.
2 Related Work
Pseudo-labeling, where a trained model generates labels for unlabeled data, has become a cornerstone of semi-supervised learning [15]. The Noisy Student framework [33] demonstrated that iteratively training larger student models on pseudo-labeled data with noise injection can surpass teacher performance, establishing a paradigm for leveraging unlabeled data at scale. FixMatch [28] combined consistency regularization with pseudo-labeling using confidence thresholds, while Mean Teacher [30] introduced exponential moving averages of model weights for stable pseudo-label generation. Meta Pseudo Labels [25] further improved teacher–student training by jointly optimizing the teacher based on student feedback. Within MIR, pseudo-labeling has been successfully applied to various tasks including piano transcription [29], music tagging [14], and speech recognition [18]. For ACR specifically, Bortolozzo et al. [1] adapted Noisy Student for rare chord recognition, employing confidence filtering and iterative teacher–student training to address class imbalance. Li et al. [16] applied contrastive learning to learn chord representations that transfer across datasets. However, these approaches typically require ground-truth labels from the outset and train on mixed pseudo-label and ground-truth signals within a single pipeline. In contrast, we investigate whether separate incremental learning stages can achieve comparable or superior performance by first training on pseudo-labels alone, then adapting to labeled data.
Transferring knowledge via temperature-scaled soft targets, or knowledge distillation (KD) [12], has been shown to provide richer supervision than hard labels. Beyond its original model compression goal, recent work has revealed KD’s regularization properties: soft labels reduce variance at the cost of introducing teacher bias [11], and Yuan et al. [34] studied conditions under which biased soft labels can still improve student generalization. Chen et al. [3] and Mansourian et al. [19] further demonstrated that KD acts as a regularizer against overfitting to noisy annotations. In continual learning, KD has been used to preserve prior knowledge while adapting to new data, mitigating catastrophic forgetting [17, 7]. Learning without Forgetting (LwF) [17] pioneered using distillation from the model’s own previous state to retain old knowledge when learning new tasks. Within MIR, KD has primarily served model compression, training smaller students to match larger teachers for efficient deployment [14]. We apply KD differently: as a regularization mechanism during continual learning that anchors the student to the teacher’s generalized representations. This preserves pseudo-label knowledge when ground-truth labels conflict with teacher predictions, while permitting adaptation when they align. Consequently, KD enables the student to refine decision boundaries without catastrophic forgetting.
Continual learning addresses the challenge of learning from non-stationary data streams without forgetting previously acquired knowledge [31]. In the MIR domain, this challenge is particularly relevant given the ongoing creation of new music and evolving annotation standards. While task-incremental and class-incremental scenarios have received significant attention, data-incremental continual learning remains underexplored for ACR. In this setting, the task and label space remain unchanged but new data arrives.
3 Methodology
3.1 Problem Formulation
Let $\mathcal{D}_{l}=\{(x_{i},y_{i})\}_{i=1}^{N_{l}}$ denote a small labeled dataset where $x_{i}\in\mathbb{R}^{T\times F}$ represents time–frequency features (e.g., Constant-Q Transform) with $T$ frames and $F$ frequency bins, and $y_{i}\in\{1,2,\ldots,C\}^{T}$ are frame-wise chord labels over a vocabulary of size $C$. Let $\mathcal{D}_{u}=\{x_{j}\}_{j=1}^{N_{u}}$ represent large-scale unlabeled datasets where $N_{u}\gg N_{l}$. Our objective is to train an effective student model $f_{s}$ by leveraging pseudo-labels generated from $\mathcal{D}_{u}$. Consequently, this mitigates reliance on expensive manual annotations while maintaining competitive performance.
3.2 Training Method
We employ a pre-trained teacher model $f_{t}:\mathbb{R}^{T\times F}\rightarrow\mathbb{R}^{T\times C}$ trained on $\mathcal{D}_{l}$ to generate pseudo-labels for unlabeled data. Any open-weight ACR model can serve as the teacher. In our experiments, we use the BTC model of Park et al. [23] as the teacher model, and use the standard vocabulary size $C=170$ for all training settings. For each unlabeled sequence $x_{u}\in\mathcal{D}_{u}$, pseudo-labels are generated via frame-wise argmax over teacher outputs:
$\hat{y}_{u}=\arg\max_{c}\left[f_{t}(x_{u})\right]_{c},\qquad\forall\,\text{frames in }x_{u}$ (1)
To preserve temporal coherence, we keep sequence boundaries intact during teacher inference (e.g., padding only at segment ends) and convert 100% of frames to pseudo-labels without confidence filtering, yielding a pseudo-labeled dataset:
$\mathcal{D}_{u}^{(p)}=\{(x_{j},\hat{y}_{j})\}_{j=1}^{N_{u}}$ (2)
Our method uses $n$ complementary unlabeled datasets where $\mathcal{D}_{u}=\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\cdots\cup\mathcal{D}_{n}$. The specific datasets used in our experiments are described in Section 4.1.
Figure 1: Duration-weighted chord root distribution across pseudo-labeled datasets. The dashed line indicates uniform distribution (8.33%). Pitch classes are well-represented with 98.4% uniformity.
Figure 2: Constant-Q Transform (CQT) comparison revealing pitch-shifting artifacts. Top row: original and pitch-shifted (5-5 semitones) spectrograms. Bottom row: artifact intensity maps for $\pm 5$ semitones, computed by realigning shifted CQT bins to compensate for the intended frequency shift.
3.2.1 Data Augmentation vs. Natural Chord Root Coverage
[…]
We observe that large-scale pseudo-labeling provides an alternative that eliminates pitch-shifting. We analyzed chord root distributions across 101,575 pseudo-labeled tracks ($\sim$}1,072 hours) from FMA, DALI, and MAESTRO. As shown in Figure 1, all 12 pitch classes are well-represented with a coefficient of variation (CV) of 0.28, substantially lower than typical labeled datasets. The Shannon entropy reaches 3.53 bits (the maximum being 3.58 bits), corresponding to 98.4% uniformity. This natural coverage arises because diverse unlabeled corpora span multiple genres, artists, and production contexts, each contributing different key preferences that aggregate to near-uniform distribution. Consequently, pseudo-label pretraining exposes the model to patterns in all keys without requiring pitch-shifting augmentation. This avoids computational overhead and potential audio artifacts from signal manipulation.
Figure 3: Illustration of the proposed two-stage training pipeline. Details of audio data are mentioned in 4.1. Note: The resulting Student Model CL of stage 2 can be continually trained when additional labeled data is available.
3.2.2 The Pipeline of Training
Figure 3 illustrates the complete two-stage training pipeline. In Stage 1, unlabeled data is first preprocessed into Constant-Q Transform spectrograms following the procedure described in Section 4.1. A pre-trained model is used as a teacher to infer frame-wise pseudo-labels (PL) on the unlabeled audio for each track, producing paired data $(Spectrogram,PL)$. A student model is then trained on these pseudo-labeled pairs until convergence. Optionally, knowledge distillation (KD) from the teacher’s soft targets can be applied during this stage to accelerate convergence; this optional path is depicted as a dashed line in the figure. We denote the resulting model after the first stage of training as the Student model PL. Stage 2 begins when the labeled data becomes available. Labeled audio undergoes the same preprocessing, yielding paired data $(spectrogram,GT)$, where GT denotes ground-truth chord annotations. In this stage, Student Model PL serves as the weight initialization for a new model called Student Model CL. Then the Student Model CL serves as the initialization for continual learning as additional labeled datasets are acquired. The selective KD information, mentioned in Section 3.2.4, from the original pre-trained teacher is incorporated throughout the training on Student Model CL. KD acts as regularization, anchoring the student to the teacher’s distributional knowledge when ground-truth labels conflict with teacher predictions, while still allowing adaptation when they agree. In both stages, the training loss is a weighted sum of a classification term and a KD term as follows:
$\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{KD}}+(1-\alpha)\mathcal{L}_{C}$ (3)
where $\alpha\in[0,1]$ controls the balance between teacher regularization and label supervision, and $\mathcal{L}_{C}$ denotes the classification loss. Specifically:
$\mathcal{L}_{C}=\begin{cases}\mathcal{L}_{PL}=\frac{1}{|\mathcal{D}_{u}^{(p)}|}\sum_{(x,\hat{y})\in\mathcal{D}_{u}^{(p)}}\ell_{CE}(f_{s}(x),\hat{y})&\text{if Stage 1}\\[7.74997pt] \mathcal{L}_{CE}=\frac{1}{|\mathcal{D}_{l}|}\sum_{(x,y)\in\mathcal{D}_{l}}\ell_{CE}(f_{s}(x),y)&\text{if Stage 2}\end{cases}$ (4)
where $\mathcal{D}_{u}^{(p)}$ denotes the pseudo-labeled unlabeled data and $\mathcal{D}_{l}$ denotes the ground-truth labeled data. Setting $\alpha{=}0$ reduces Eq. (3) to pure classification training without KD. The KD loss is defined as:
$\mathcal{L}_{\text{KD}}=\tau^{2}\cdot D_{\text{KL}}\!\left(\sigma\!\left(\frac{\mathbf{z}_{t}}{\tau}\right)\Big\|\sigma\!\left(\frac{\mathbf{z}_{s}}{\tau}\right)\right)$ (5)
where $\sigma(\cdot)$ denotes the softmax function, $\tau>0$ is the temperature parameter that controls the smoothness of probability distributions, $D_{\text{KL}}$ is the Kullback-Leibler divergence, and $\mathbf{z}_{s},\mathbf{z}_{t}\in\mathbb{R}^{C}$ are the student and teacher logits, respectively.
3.2.3 KD as Regularization
We denote the temperature-softened probability distributions as $\mathbf{p}^{(s)}=\sigma(\mathbf{z}_{s}/\tau)$ and $\mathbf{p}^{(t)}=\sigma(\mathbf{z}_{t}/\tau)$, with $p^{(s)}_{k}$ and $p^{(t)}_{k}$ denoting the $k$-th class probability. The $\tau^{2}$ scaling ensures gradient magnitudes remain comparable to cross-entropy [12]. The KD gradient with respect to student logits is:
$\nabla_{\mathbf{z}_{s}}\mathcal{L}_{\text{KD}}=\tau\,(\mathbf{p}^{(s)}-\mathbf{p}^{(t)}).$ (6)
This gradient “pulls” student predictions toward the teacher’s distribution. Combining with the classification term from Eq. (3):
$\frac{\partial\mathcal{L}_{\text{total}}}{\partial z_{s,k}}=(1-\alpha)\frac{\partial\mathcal{L}_{C}}{\partial z_{s,k}}+\alpha\tau\big(p^{(s)}_{k}-p^{(t)}_{k}\big)$. (7)
The second term acts as a regularizer that anchors student predictions to the teacher’s distribution. When ground-truth labels conflict with teacher predictions (e.g., due to annotation noise or misalignment), this regularization prevents overfitting to erroneous labels. Conversely, when labels align with teacher predictions, the regularization does not impede adaptation. This property is particularly beneficial for ACR, where annotation inconsistencies are common due to subjective harmonic interpretation [9, 24].
3.2.4 Selective KD
To further improve robustness, we apply selective KD that filters the teacher signal based on prediction confidence. Let $c=\max_{i}p_{i}^{(t)}$ denote the teacher’s maximum softmax probability. We define an asymmetric weighting function, $w(c)\in[0,1]$ and:
$w(c)=\begin{cases}0&\text{if }c<\theta_{min}\\ 1&\text{if }\theta_{min}\leq c\leq\theta_{max}\\ 1-K\cdot\frac{c-\theta_{max}}{1-\theta_{max}}&\text{if }c>\theta_{max}\end{cases}$ (8)
where $\theta_{min}$ filters out uninformative low-confidence samples, $\theta_{max}$ down-weights overconfident predictions that may bias toward majority classes, and factor $K\leq 1$ controls the overconfident predictions rate. The weighted KD loss becomes $\mathcal{L}_{KD}^{sel}=\frac{1}{N}\sum_{i}w(c_{i})\cdot\ell_{KD}(x_{i})$. Samples near decision boundaries (moderate confidence) contain valuable uncertainty information, so we preserve full weight in the informative range $\theta_{min},\theta_{max}]$. We apply selective KD with $\theta_{min}=0.1,\theta_{max}=0.9,K=0.8$ throughout all Stage 2 continual learning experiments to stabilize training by reducing gradient variance from extreme-confidence samples.
Figure 4: Experimental Dual Encoder Architecture (2E1D): The model consists of separate temporal and frequency encoders that process CQT features independently before fusion for chord classification.
3.2.5 Experimental Models
BTC [23] serves as both the pseudo-labeling teacher and the baseline for self-distillation experiments. It uses a CNN front-end for local spectral feature extraction followed by a bi-directional transformer for temporal modeling, representing a deeper architecture with stacked convolutional and attention layers.
For cross-architecture validation, we introduce a new compact dual-encoder architecture (2E1D) for chord recognition that is purely transformer-based [32] (without any CNN layer). As shown in Figure 4, 2E1D adopts a wider design: (1) a frequency encoder that groups CQT bins into spectral clusters and applies self-attention to learn harmonic relationships across frequency bands; (2) a temporal encoder that processes full-band features to model chord progression patterns over time; and (3) a cross-attention fusion block that combines the two streams for final chord classification. At inference time, we apply a temporal smoothing pipeline over output logits to reduce frame-level prediction jitter. Each class logit channel is convolved with a normalized 1D Gaussian kernel $g[n]=\exp(-n^{2}/2\sigma^{2})/Z$, where $\sigma=k/6$ follows the three-sigma rule for a kernel of width $k$, with replicate padding at segment boundaries. We further process the input using overlapping sliding windows with a stride of $s=\lfloor T(1-r)\rfloor$, where $T$ is the segment length and $r$ is the overlap ratio, accumulating per-class votes across all windows before taking the frame-wise argmax. The use of a wider architecture is to trade depth for parallel processing capacity: the 2E1D model generally processes faster than BTC. We investigate whether attention alone can capture both local spectral patterns and global temporal dependencies without CNN inductive biases [24, 6]. As shown in Section 5, the wider 2E1D architecture is more susceptible to degradation when adapting to noisy labels compared to the deeper architecture of BTC. Thus, the 2E1D model requires stronger KD regularization to maintain stability.  [^1] Models: https://github.com/ptnghia-j/ChordMini
4 Experiments
4.1 Datasets and Preprocessing
Unlabeled Datasets. We use three large-scale unlabeled datasets for pseudo-label generation, totaling over 1,000 hours of audio. The Free Music Archive [4] ($\mathcal{D}_{FMA}$) provides over 100,000 short-form tracks ($\sim$30s) with extensive genre diversity. The MAESTRO dataset [10] ($\mathcal{D}_{MAESTRO}$) contributes 200+ hours of high-quality piano recordings with precise MIDI alignment. The DALI dataset [22] ($\mathcal{D}_{DALI}$) supplies over 5,000 full-length music tracks.
Labeled Datasets. We aggregate annotations from the Isophonics dataset [20], the McGill Billboard corpus [2], RWC Pop [8], and USPop datasets [21], collecting a fixed subset of 600 songs. The dataset is then split with a ratio of 7:1:2 into train/validation/test sets (420/60/120 songs). The “50%” and “full” conditions in our continual learning experiments refer to using 210 and 420 training songs from this subset, respectively.
Clean vs. Noisy Annotations. To evaluate KD’s robustness to annotation noise, we prepare two versions of the labeled data: (1) clean labels with careful manual alignment, and (2) noisy labels sourced online without alignment correction, primarily affecting non-chord (“N”) label boundaries. This setup enables an ablation study of KD’s regularization effect under different noise conditions.
Preprocessing. All audio undergoes identical preprocessing. The Constant-Q Transform (CQT) features [27] are extracted with $F=144$ frequency bins, 24 bins per octave, and a hop length of $h=2048$ samples, yielding a temporal resolution of $\Delta t\approx 93$ms at $f_{s}$= 22.05 kHz. Features undergo z-score normalization using teacher model statistics ($\mu_{t}$, $\sigma_{t}$) to maintain identical input distributions.
4.2 Training Configuration
In the first stage, pseudo-labeling training employs the AdamW optimizer with a learning rate $10^{-4}$, a batch size of 256, and a sequence length of 108 frames ($\sim$10 seconds). The learning rate schedule incorporates warmup over 10 epochs to $3\cdot 10^{-4}$, followed by cosine annealing decay. Early stopping monitors validation accuracy with a patience of 10 epochs. We reserve 10% of the pseudo-labeled data for validation and 10% as a held-out test set. In the second stage, continual learning, we use a reduced learning rate of $10^{-5}$ with decay upon observing a validation plateau. KD weight $\alpha{=}0.3$ balances adaptation and regularization. The temperature $\tau{=}3.0$ is empirically selected as the optimal value in all settings. Training continues until early stopping triggers. No data augmentation is applied in our two-stage pipeline (Section 3.2.1).
For the supervised learning (SL) baseline, both BTC and 2E1D are trained from scratch on the full labeled training set (420 songs) using the AdamW optimizer with a learning rate of $10^{-4}$, a batch size of 256. Learning rate decays with rate 0.95 when validation accuracy did not improve. Pitch-shifting augmentation via the Rubber Band library [5] transposes both audio and labels from 5-5 to +6+6 semitones.
4.3 Evaluation Metrics
We employ standard metrics from mir_eval [26]: Root, Thirds, Triads, Sevenths, Tetrads, Majmin, and MIREX (overall chord recognition). Frame-wise accuracy (Acc), precision (Prec), and recall (Rec) are computed using standard True Positive, False Positive, and False Negative counts. Additionally, we report Chord Symbol Recall (CSR) and Weighted Chord Symbol Recall (WCSR) following Bortolozzo et al. [1]:
$CSR_{i,j}=\frac{|S_{i}\cap S_{j}|}{|S_{j}|},\qquad WCSR_{i}=\frac{\sum_{j}T_{j}\cdot CSR_{i,j}}{\sum_{j}T_{j}}$ (9)
where $S_{i}$ denotes predicted chord labels, $S_{j}$ represents ground-truth labels, and $T_{j}$ is the duration of the $j$-th track. The Average Chord Quality Accuracy (ACQA), with chord set $C$, is calculated as:
$ACQA=\frac{\sum_{C_{i}\in C}WCSR_{C_{i}}}{|C|}$ (10)
WCSR weights accuracy by class distribution, while ACQA gives equal weight to all chord qualities. This makes ACQA more sensitive to rare chord performance.
We also report segmentation metrics from mir_eval [26]. Let $\mathcal{I}_{ref}$ and $\mathcal{I}_{est}$ denote the sets of reference and estimated chord segment intervals, respectively. Over-segmentation (Over) measures how often predicted boundaries subdivide reference segments:
$\text{Over}=1-\frac{\sum_{I\in\mathcal{I}_{ref}}\max_{J\in\mathcal{I}_{est}}|I\cap J|}{\sum_{I\in\mathcal{I}_{ref}}|I|}$ (11)
Under-segmentation (Under) measures how often reference boundaries subdivide predicted segments:
$\text{Under}=1-\frac{\sum_{J\in\mathcal{I}_{est}}\max_{I\in\mathcal{I}_{ref}}|I\cap J|}{\sum_{J\in\mathcal{I}_{est}}|J|}$ (12)
Let $N$ denote the number of songs in the evaluation set. The overall segmentation score (Seg) uses the maximum directional Hamming distance, macro-averaged across songs:
$\text{Seg}=\frac{1}{N}\sum_{i=1}^{N}\bigl(1-\max\{\text{Over}_{i},\,\text{Under}_{i}\}\bigr)$ (13)
mir_eval Metrics Segmentation Frame-wise vs. teacher
Model Training Data Root Thirds Triads 7ths Tetrads Majmin MIREX Over Under Seg Acc Prec Rec F1
Pre-trained teacher model (reference)
BTC [23] Original weights 82.28 78.81 77.07 67.00 64.25 79.31 78.71 81.74 90.34 80.43
Our two student models trained only with pseudo-labels (generated by the BTC teacher from unlabeled data)
2E1D (2.2M) FMA (short-form) 77.28 73.70 72.15 61.78 58.94 74.49 73.68 83.55 83.96 78.03 57.27 34.43 20.29 23.51
DALI (long-form) 74.32 65.53 63.73 52.68 50.27 65.91 65.37 77.09 85.92 74.00 78.25 48.06 37.41 39.52
MAESTRO+DALI 79.50 75.80 74.11 63.18 60.28 76.53 75.50 83.14 84.85 78.53 79.44 52.23 41.95 44.50
All (FMA+M+D) 80.37 76.63 74.91 64.37 61.50 77.29 76.35 84.34 85.63 80.15 80.15 57.01 45.20 47.49
All + KD ($\alpha{=}0.3$) 80.17 76.36 74.68 63.93 61.06 77.09 76.08 84.23 85.27 79.69 81.23 58.14 46.33 48.62
BTC (3.03M) FMA (short-form) 78.85 74.76 73.12 62.88 60.09 75.37 75.07 80.92 87.57 78.44 61.12 36.53 24.11 26.37
DALI (long-form) 81.01 77.24 75.35 65.46 62.59 77.66 77.14 79.50 90.60 78.51 85.88 49.23 50.37 47.32
MAESTRO+DALI 80.65 76.37 74.61 64.19 61.35 76.93 76.69 77.42 90.67 76.66 88.01 58.84 64.29 59.16
All (FMA+M+D) 81.54 77.86 76.08 66.29 63.54 78.29 77.84 80.97 90.15 79.76 89.14 62.54 63.89 59.34
All + KD ($\alpha{=}0.3$) 81.23 77.75 75.94 66.00 63.22 78.15 77.82 80.81 90.25 79.60 89.47 63.21 64.05 60.12
Table 1: Pseudo-labeling results across dataset configurations. The mir_eval metrics are computed against ground-truth test set, while frame-wise accuracy, precision, recall, and F1 measure agreement with the teacher’s predictions. FMA contains short-form clips ($\sim$30s), while DALI and MAESTRO provide long-form full tracks. The pre-trained BTC teacher [23] serves as the upper bound. Best results for BTC and 2E1D in the table are highlighted in blue and green, respectively.
5 Results
In this section, we present results across the various metrics mentioned above in our two-stage training method, including an ablation study on the KD regularization against noisy labels. All reported metrics are evaluated on the held-out test set (120 songs) from the clean labeled dataset described in Section 4.1.
5.1 Stage 1: Training with pseudo-labels
Increasing unlabeled data diversity consistently improves performance across mir_eval metric hierarchy and segmentation quality (Table 1). Training on all three datasets (FMA, MAESTRO, DALI) yields the strongest results for both architectures, indicating that pseudo-label pretraining benefits from coverage of varied genres and recording conditions. Long-form datasets consisting of full-length tracks (DALI, MAESTRO) produce more stable pseudo-labels than short-form clips (FMA): DALI achieves higher frame-wise agreement with the teacher than FMA. Longer musical contexts provide consistent harmonic progressions, reducing boundary jitter and spurious chord predictions. However, FMA achieves better mir_eval results than DALI due to its larger dataset size. The best BTC student reaches over 98% of teacher performance on ground-truth evaluation across all mir_eval, while the purely transformer-based 2E1D achieves 96-97% of the results from the teacher model, demonstrating cross-architecture knowledge transfer. When all training targets are pseudo-labels, adding KD aligns the student to the teacher’s soft-label distribution better, resulting in improvements across frame-wise metrics despite a decrease in mir_eval metrics. In addition, KD accelerates optimization, as all pseudo-label trainings run with KD converge in 10 epochs versus 20–40 for other settings; this comes with a modest training-time memory overhead to store and backpropagate full soft targets  [^2] A chord recognition application using the model by our training method is available at: https://github.com/ptnghia-j/ChordMiniApp.
5.2 Ablation: KD as Regularization
To understand KD’s role as a stabilizer under label noise, we evaluate continual learning with misaligned labels at varying $\alpha$ values (Table 2). Without KD ($\alpha{=}0$), both architectures suffer severe and consistent degradation across all metrics: BTC drops substantially throughout the recognition hierarchy, while the wider 2E1D collapses far more drastically to near-failure levels. The dual attention architecture is more susceptible to noisy label overfitting than the deeper CNN-transformer BTC. KD acts as an anchor, preventing noisy fine-tuning from destroying a good pseudo-label initialization. KD adds robustness without impeding adaptation. BTC peaks at $\alpha{=}0.3$ with a broad, consistent recovery across all metrics relative to the no-KD baseline, while 2E1D requires stronger regularization ($\alpha{=}0.5$) to stabilize. Figure 5 visualizes training dynamics. Without KD ($\alpha{=}0$), validation loss rises as the model overfits to noisy labels, whereas when $\alpha>0$}, KD anchors the student to the teacher’s distribution. This supports KD as an effective safeguard that preserves pseudo-label knowledge during adaptation.
Figure 5: Evaluation loss of BTC model during continual training with different KD weights. Higher $\alpha$ values provide stronger regularization, mitigating performance degradation from noisy labels.
Model $\alpha$ Root Thirds Triads 7ths Tetrads Majmin MIREX Seg
2E1D 0 52.47 50.70 49.88 42.52 41.00 51.12 50.99 66.72
0.1 55.85 54.16 53.41 47.68 45.78 54.88 54.07 67.70
0.3 66.05 63.89 62.79 55.78 53.39 64.70 63.87 73.32
0.5 74.21 71.25 69.78 61.20 58.35 72.09 71.54 76.84
BTC 0 71.06 68.05 66.73 58.11 55.66 68.69 68.20 76.05
0.1 73.88 70.86 69.43 60.91 58.25 71.52 70.96 77.59
0.3 77.34 73.95 72.44 62.97 60.25 74.60 74.29 78.61
0.4 76.95 72.88 71.30 61.30 58.57 73.59 73.18 77.98
Table 2: KD regularization effect during continual training with misaligned labels. The KD weight $\alpha$ controls the balance between teacher soft targets and ground-truth hard labels. BTC peaks at $\alpha{=}0.3$, while 2E1D requires stronger regularization. Best results for BTC and 2E1D are highlighted in blue and green, respectively.
mir_eval Metrics Per-Chord Quality Accuracy Aggregate
Method Root Thirds Triads 7ths Tetrads Majmin MIREX Seg Maj Min Dom7 Maj7 Min7 Dim Dim7 Aug Sus WCSR ACQA
Prior Work (Noisy Student Framework)
Bortolozzo et al. [1] 55.5 54.4 51.0 12.0 47.5 30.4 43.3 46.8 24.5
Li et al. [16] 78.05 73.56 70.71 61.11 56.84 74.64 74.26
Traditional supervised learning (SL) baseline (trained on labeled data (full labels) with data augmentation
2E1D (SL) (2.2M) 79.39 76.14 74.53 64.53 61.67 76.89 76.24 78.50 91.7 62.1 23.9 25.2 44.4 3.5 0.0 0.0 8.0 68.7 24.4
BTC (SL) (3.03M) 81.52 78.00 76.12 65.93 63.44 78.11 77.79 78.38 88.7 63.1 36.7 28.2 52.8 23.3 0.0 5.9 22.9 68.6 39.0
Our student models resulted in Stage 2 with Data-Incremental Continual Learning (with clean labels, $\alpha{=}0.3$)
2E1D (CL) (50%) 81.02 77.54 75.76 66.39 63.68 77.93 77.57 80.23 90.9 63.0 33.3 29.5 49.0 12.7 46.7 5.2 16.6 70.4 31.2
2E1D (CL) (full) 81.55 78.52 76.85 68.19 65.49 78.98 78.52 80.73 90.0 68.9 39.4 35.3 51.9 14.8 46.4 8.4 18.8 71.9 33.7
BTC (CL) (50%) 82.14 78.58 76.67 66.77 64.28 78.62 78.87 80.17 84.5 61.2 50.2 42.1 51.6 18.3 47.4 14.8 22.3 68.7 47.1
BTC (CL) (full) 83.03 80.17 78.33 69.38 67.00 80.24 80.16 80.71 87.1 65.6 51.6 42.3 55.7 23.5 45.6 18.5 25.5 71.6 49.5
Table 3: Data-incremental continual learning results with per-chord quality accuracy. After pseudo-labeling, we fine-tune on ground-truth labels with KD ($\alpha{=}0.3$). Our student BTC (CL) with full labels surpasses the teacher BTC shown in Table 1 across all mir_eval metrics. Continual learning significantly improves rare chord recognition (Dim, Dim7, Aug) compared to the supervised baseline. Best results for BTC and 2E1D are highlighted in blue and green, respectively. We report only the metrics available in the original papers with our test set; cells not reported are left empty.
5.3 Stage 2: Continually training with GT + KD
BTC with full labels surpasses both the teacher (Table 3) and the supervised learning (SL) baseline consistently across all seven mir_eval metrics. The 2E1D model similarly achieves broad improvements over its SL baseline. Among all metrics, 7ths and Tetrads benefit most from the combined effect of pseudo-label pretraining and ground-truth fine-tuning. Stage 1 supplies broad but biased coverage from diverse unlabeled audio, and Stage 2 corrects this bias using ground-truth. Crucially, KD does not prevent improvement when labels are clean (Table 3): models still surpass the teacher, indicating that KD selectively mitigates noise while preserving adaptation capacity. The complete two-stage pipeline demonstrates two key insights: (1) pseudo-label pretraining reduces labeled data requirements, as with a minimal additional labeled dataset, BTC can surpass the teacher; and (2) gains are disproportionately concentrated on rare chord qualities, consistent with our scarcity motivation. The ACQA metric, which weights all chord qualities equally, reveals disproportionate gains: BTC improves by 10.5% from 39.0% to 49.5%, compared to only 3.0% improvement for distribution-weighted WCSR. This contrast strongly supports the observation that pseudo-label pretraining provides decision boundaries for rare chords that labeled datasets alone cannot establish. Figure 6 visualizes this effect. Major (45.5%) and Minor (22.3%) chords dominate the training distribution, while rare qualities (Dim: 1.8%, Dim7: 0.4%, Aug: 0.3%) constitute less than 3% combined. The SL baseline achieves 0% on Dim7 due to insufficient examples, while our pipeline leverages teacher pretraining to achieve 45.6%, a shift from failure to reasonable recognition.
Comparison to Prior Work. Table 3 demonstrates that pseudo-label quality is the critical determinant of semi-supervised learning performance. Prior work [1, 16] relies on supervised models trained on limited labeled data to generate pseudo-labels. This drastically constrains label quality and results in weaker model performance. In contrast, training with high-quality pseudo-labels generated by a better teacher model provides better initialized-weights, thereby creating a stronger starting point for model training and adaptation.
Figure 6: Chord quality distribution (bars) and recognition accuracy (lines) of models trained with traditional supervised learning approach (BTC (SL), 2E1D (SL)) versus our approach (BTC (CL), 2E1D (CL)).
6 Conclusion
Since model weights are often more readily available than proprietary training data, we proposed a practical training strategy for the ACR problem that leverages open-weight pretrained models when high-quality labels are scarce. We show that students trained solely on pseudo-labels can approach teacher-level performance across seven mir_eval metrics. We further demonstrate that continual learning can improve performance without catastrophic forgetting when the teacher provides sufficiently general representations. Under our pipeline, the student ultimately surpasses the teacher, especially on rare chord qualities (e.g., Dim, Dim7, Aug). Knowledge distillation improves robustness to noisy labels while preserving adaptability when ground-truth annotations are clean. We also find that the wider 2E1D architecture requires stronger KD regularization than the deeper BTC, underscoring how architectural choices influence continual-learning stability. A key limitation is reliance on teacher quality: biased or weakly generalizable teacher representations can transfer these shortcomings to the student. Future work will explore stronger teachers, ensembles of multiple teachers, scaling to additional unlabeled corpora, and extending the framework to related MIR tasks such as beat tracking and key estimation.